OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (−2, 11)?

OpenStudy (anonymous):

first find the slop by using \[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]

OpenStudy (anonymous):

you'll get \[\frac{ 5-11 }{ 0+-2 }\] which is equal to -3

OpenStudy (anonymous):

the next step is to plug in your values to the point-slope equation \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

you should get y-5=-3(x-0) here i used the point (5,0) but the other one works just fine

OpenStudy (anonymous):

this should be your final answer if it asks for point-slope form. if it asks for y-intercept form, you can continue to solve using algebra (distribute the -3, then add 5 to both sides)

OpenStudy (anonymous):

sorry i meant (0,5) not (5,0)